OpenStudy (ucreategames):

in triangle abc, the measure of angle b is two times the measure of angle a. the measure of angle c is twenty degrees more than angle b. find the measure of each angle.

OpenStudy (anonymous):

A=32,B=64,C=84

OpenStudy (ucreategames):

thnx

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

a=b/2 c=b+20 so b+(b+20)+(b/2)=180 2.5b+20=180 2.5b=160 b=64 a=32 c=84

OpenStudy (ucreategames):

thnx